OpenStudy (anonymous):

evaluate the limit by first recognizing the sum as a rieman sum on the interval 0 to 1. lim as n approaches infinity, (1/sqrt(n))((1/sqrt(n+2)+(1/sqrt(n+3)...+1/sqrt(n+n)

OpenStudy (anonymous):

\[\int_0^1\frac{1}{\sqrt{x+1}}dx\] is my guess

OpenStudy (anonymous):

i am assuming you are missing a plus sign at the beginning

OpenStudy (anonymous):

no its listed a a reimann sums and for the other problems we had to find a F(x)

OpenStudy (anonymous):

(1/sqrt(n)) is delta x

OpenStudy (anonymous):

ok so if \[\Delta x=\frac{1}{\sqrt{n}}\] then that means \[f(\frac{1}{\sqrt{n}})=\frac{1}{\sqrt{n+1}}\] and \[f(\frac{2}{\sqrt{n}})=\frac{1}{\sqrt{n+2}}\] ektc

OpenStudy (anonymous):

what the heck is this integral???

OpenStudy (anonymous):

or am i completely on the wrong track...

OpenStudy (anonymous):

seems like i was off by 1 \[f(\frac{1}{\sqrt{n}})=\frac{1}{\sqrt{n+2}}\] i must be doing something wrong

OpenStudy (zarkon):

is this the problem \[\frac{1}{\sqrt{n}}\left[\sum_{i=2}^{n}\frac{1}{\sqrt{n+i}}\right]\]

OpenStudy (anonymous):

i think so yes, and the question is what integral is this

OpenStudy (anonymous):

on the interval [0,1] and letting n go to infinity maybe i am supposed to be thinking \[\Delta x=\frac{1}{\sqrt{x}}\] and then find x? i am totally at sea here

OpenStudy (zarkon):

\[=\frac{\sqrt{n}}{n}\left[\sum_{i=2}^{n}\frac{1}{\sqrt{n+i}}\right]\] \[=\frac{1}{n}\left[\sum_{i=2}^{n}\frac{\sqrt{n}}{\sqrt{n+i}}\right]\] \[=\frac{1}{n}\left[\sum_{i=2}^{n}\frac{1}{\frac{\sqrt{n+i}}{\sqrt{n}}}\right]\] \[=\frac{1}{n}\left[\sum_{i=2}^{n}\frac{1}{\sqrt{1+\frac{i}{n}}}\right]\]

OpenStudy (anonymous):

wow so my initial guess was correct?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

second luckiest guess here. earlier this summer someone posted a series and said "what is this?" as a challenge and i wrote \[\frac{e^2}{4}\] as a joke and it was right.

OpenStudy (zarkon):

nice

OpenStudy (anonymous):

i like to think of it as "intuition" but it really was random